OpenStudy (anonymous):

John can paint a room in 3 hours, Mark can paint a room in 12 hours. They want to paint a room together to save time. How long would it take?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

In 1 hr, John will paint 1/3 of room In 1 hr, Mark will paint 1/12 of room Together in 1 hr, they will paint 1/3 + 1/12 = 5/12 of room So together they will finish the work in 12/5 hrs = 2 hr 24 min